OpenStudy (anonymous):

Hi , It doesnt seem like this can be solved to x : (3x+2)(4x+10)

OpenStudy (anonymous):

are these equal to zero?

OpenStudy (anonymous):

(3x+2)(4x+10) =0 ?

OpenStudy (anonymous):

I end up with the following awnser 6x2 + 38x + 20 = 0 so nothing in c will multiply to the sum of b

OpenStudy (anonymous):

In the book it is not equal to 0 it says evaluate the following (3x+2)(4x+10)

OpenStudy (anonymous):

(3x+2)(4x+10) =12x^2 + 8x+30x+20 =12x^2 +38x +20 thats all you can do to it unless you want to find its root or zeros

OpenStudy (anonymous):

ok o just simplified it more. wont affect any marks ?

OpenStudy (anonymous):

woops i wrote it wrong 6x2+19x+10 = 0

OpenStudy (anonymous):

so thats the most simplified awnser.

OpenStudy (anonymous):

that ok, so if you want to equate them to zero, either you use the quadratic formula or the one already factored earlier w/c result x=-2/3,-5/2

OpenStudy (anonymous):

ok so if i cant solve it ill have to use x = -b +_ square root b 2 - 4ac / 2a

OpenStudy (anonymous):

no, only if the problem says determine its zero or roots by equating them to zero, also if you cant factor them anymore, then you can also use the quadratic formula .... try to solve it using the formula lets see what the zeroes